OpenStudy (anonymous):

A band recently sold 5000 tickets to a concert. If there are 2 types of tickets, regular and children's and there were 600 more children's ticket sold than regular tickets, how many children's tickets were sold? HOW TO SOLVE?

OpenStudy (anonymous):

@jim_thompson5910 @mathmate

OpenStudy (danjs):

Set up two Equations with 2 unknown values. Total Tickets Sold = #Children + # Regular and #Regular + 600 = #Children

OpenStudy (danjs):

Total Tickets Sold is 5000 Solve for #Regular and # Children

OpenStudy (danjs):

Let #Regular = R and #Children = C, the two equations become 5000 = R + C C = R + 600 Solve for R and C.

OpenStudy (anonymous):

Im confused

OpenStudy (mathmate):

 Look at the diagram above. R=# of regular tickets. Children outnumber regular by 600. All of them add up to 5000 So set up R + (R+600) = 5000 Solve for R (# of regular tickets)